the probability of winning something on a single play at a slot machine is 0.13. after 5 plays on the slot machine, what is the probability of winning at least once
Answer #1

X ~ binomial (n,p)

Where n = 5 , p = 0.13

P(X) = nCx px ( 1 - p)n-x

P(X >=1 ) = 1 - P(X = 0)

= 1 - 5C0 0.130 ( 1 - 0.13)5

= 0.3724